A hospital has 100 beds in 50 semi private rooms. Average occupancy over the past several years has been 70%. Determine the number of projected encounters using the 70% occupancy factor. Determine the number of weighted encounters under the assumption that the assigned weight is 4.0.


1
Expert's answer
2021-02-04T14:07:52-0500

Beds = 100

Average occupancy = 70 %

Encounters "= 100 \times 365 \times 70 \% = 25550"

Weighted encounters = weightage "\times" encounters "= 25550 \times 4.0 = 102200"

Revenue budget = weighted encounters "\times" value per weighted encounter "= 102200 \times 100 = 10220000"

Revenue budget post discount factor "= 10220000 \times (100 \% -40 \%) = 6132000"